Sixty-three percent of U.S. adults oppose special taxes on junk food and soda. You randomly select 10 U.S. adults. Find the mean, variance, and standard deviation of this binomial distribution.
Answer #1

Let X be number of adults oppose special taxes on junk food and soda. Sample size ( n ) = 10

Here, X ~ Binomial ( n = 10, p = 0.63)

Mean = E(X) = np = 10 * 0.63 = 6.3

Varaince = Var(X) = np(1-p) = 10 * 0.63 * (1-0.63) = 2.331

standard deviation = SD(X) = V2.331 -1.5268

Hence,

Mean = 6.3

Variance = 2.331

Standard deviation = 1.5268